# Constrained Measurement Incompatibility from Generalised Contextuality of Steered Preparation

### v2

@@ III.2 Ontological model of a GPT and generalised noncontextuality
At this point, we recall an important result from Ref. [20] that will be used to establish our main result. We encapsulate that result in the following theorem and provide a straightforward proof of it.
@@ IV Notion of measurement incompatibility in a GPT
Given the state space $\Omega$, any measurement in the GPT formalism is represented by an affine map $M:\Omega\rightarrow\mathcal{P}(\Omega)$, where $\mathcal{P}(\Omega)$ constitutes the set of probabilities for the outcomes. Technically, a measurement $M$ on a system state $\omega\in\Omega$ produces dichotomic outcomes $m\in\{\pm 1\}$ can be represented as, $M=\{e_{\tiny M,\omega}^{\tiny m}|\sum_{m}e_{\tiny M,\omega}^{\tiny m}=1,e_{\tiny M,\omega}^{\tiny m}\geq 0,\forall\omega\in\Omega\}$. Therefore, the expected average value of $M$ can be written as
@@
Such joint effects corresponding to a pair of observables may not always exist when only perfect measurements are considered. However, for the imperfect or fuzzy measurement of a pair of observables, such a joint effect may exist if the degree of fuzziness is upper bounded by a certain critical value [47]. We call such fuzzy version of the measurements as unsharp measurement and represent the unsharpness parameter as $\eta\in[0,1]$.
@@
$N^{\eta^{*}}$-compatible GPT: A specific GPT is said to be $N^{\eta^{*}}$-compatible GPT if the necessary condition for the $N$-wise compatibility of the fuzzy measurement of $N$ arbitrary observables is $\eta\leq\eta^{*}$.
Note that for any given value of $N$, the higher the value of $\eta^{*}$, the less $N$-incompatible the GPT is. For instance, the extreme case of $\eta^{*}=1$ indicates that the GPT is $N-$wise compatible, even if the measurements are perfect. Such GPTs are deemed classical in terms of incompatibility, as those $N$ observables have simultaneous values. Following definition 3, a GPT can be said to be maximally incompatible if in such a GPT, the highest amount of unsharpness $(\eta^{\ast})$ needs to be given to make any arbitrary $N$ observable compatible. It is shown in [48] that if $\eta^{*}=\frac{1}{N}$ then all $N$ observables are compatible in any arbitrary GPT. However, it may be possible that in a specific GPT for $\eta<\frac{1}{N}$ is enough to establish $N$-wise compatibility. Therefore, a GPT is referred to as maximally incompatible if $\eta^{*}=\frac{1}{N}$ [48].
We note here that in quantum theory, the maximally incompatible pair of observables do not exist [48, 49] in finite-dimensional Hilbert space. However, if the dimension of the quantum system is infinite, one needs the unsharpness parameter $\eta=1/2$ to make an arbitrary pair of observables compatible. On the other hand, any pair of qubit observables becomes compatible if and only if the unsharpness parameter is $\eta\leq 1/\sqrt{2}$ [1]. Later, this result is generalised [12] for any pair of dichotomic observables in an arbitrary finite dimension of Hilbert space. We remark here that a very few results regarding $N$-wise compatibility in any arbitrary dimension $d$ are known. The best known [50] tight bound on the unsharpness for the compatibility of set of $N$ noisy measurements is $\eta\leq\frac{H_{d}-1}{d-1}$, where , $H_{d}=\sum_{n=1}^{d}\frac{1}{n}$ is called the $d^{th}$ harmonics. It has also been argued in [50, 21, 51] that for any dimension $d$, the above bound is necessary and sufficient for the compatibility of arbitrary number of measurements. For example, an arbitrary $N$ noisy qubit measurements becomes compatible iff $\eta\leq 1/2$ [21]. By noting the connection between compatibility of a set of measurements with quantum steering, this bound is also supported by recent results [53, 52].
Given the above perspective, we ask the more general question of what is the necessary and sufficient condition for $N$-wise compatibility for any no-signalling GPT. From the quantum mechanical perspective, such a $N$-wise compatibility condition enables us to find the critical value of $\eta$ required to make an arbitrary $N$ number of dichotomic observables compatible.
@@ V Landscape of the arguments and Results
where $\omega^{\lambda}$ is the GPT state corresponding to variable $\lambda$. Non-existence of such decomposition of the states is usually referred to as steering [54, 50] in literature. However, we do not intend to discuss the very details of steering. Instead, we use the proof for the existence of states in the form Eq. (19) as a helping tool for proving one of our main results articulated in Theorem 2, appearing later in this section.
To show that a decomposition as Eq. (19) exists for the reduced states at Bob’s wing, let us first assume that there is a one-to-one correspondence between the elements of $\Lambda$ and $a$ such that $\lambda=a_{x}$ specifies one of the remotely prepared states. The variable $\lambda$ is distributed according to the probability distribution $p(\lambda)=\langle e^{a_{x}}_{A},\omega_{A}^{u_{B}}\rangle$, where $\omega_{A}^{u_{B}}$ is Alice’s part of the state when Bob measures the unit effect $u_{B}$ on his part of the subsystem. In each run of experiment Alice sends her local state $\omega^{a_{x}}_{B}=[(e^{a_{x}}\otimes u_{B}),\omega_{AB}]$ to Bob. Furthermore, when Bob asks her to perform a specific measurement, say $x$, and announce the outcome, she simply conveys the outcome $a_{x}$ according to the probability distribution $p(a_{x}|x,\lambda)=\delta_{\lambda,a_{x}}$. Thus each of the states that Bob gets can be written as Eq. (19). The correspondence between the existence of such decomposition of remotely prepared states and the joint measurability has also been demonstrated in ref. [56, 55].
@@
Figure 1: Pictorial depiction of main results. Dark red arrows are findings of the current paper while the blur arrow indicates the result of ref. [20].
@@
$\mathcal{B}_{N}=\max_{B_{y}}\Big{(}\lvert\sum_{x=1}^{N}\sum_{y=1}^{2^{N-1}}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert\Big{)}\leq 2^{N-1}$ (29)
where $l_{y}^{x}$ is the $x^{th}$ bit sampled from $y^{th}$ string from the bit-string $l\in\{0,1\}^{n}$ having first entry 00.
@@ VI Optimal measurement incompatibility allowed in quantum theory
Consider that Alice’s measurements are described by unbiased POVMs (smeared versions of projective measurement) of the form $A_{x}\equiv\Big{\{}A_{\pm|x}\ \big{|}\ E_{\pm|x}=\frac{1}{2}(\mathbb{I}\pm\eta A_{x})\Big{\}}$, where $\eta\in[0,1]$ is the unsharpness parameter. Hence, effective observable of Alice is $\eta A_{x}$. If we assume the same unsharpness parameter $\eta$ for each of the observable $A_{x}$, we can then write the quantum value of the functional $\mathcal{B}_{N}$ as
$(\mathcal{B}_{N})_{Q}=\eta\sum_{x=1}^{N}\sum_{y=1}^{2^{N-1}}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle.$ (31)
To optimize $(\mathcal{B}_{N})_{Q}$ we use an elegant sum-of-square approach [57] as explicitly described in the following. This implies showing $(\mathcal{B}_{n})_{Q}\leq\beta_{N}$ for all possible quantum states $\rho_{AB}$ and measurement observables $A_{x}$ and $B_{y}$, where $\beta_{N}$ is the upper bound. This is equivalent to showing that there is a positive semidefinite operator $\gamma_{N}\geq 0$, that can be expressed as $\langle\gamma_{N}\rangle_{Q}=\beta_{N}-(\mathcal{B}_{n})_{Q}$. We prove this by considering a set of suitable positive operators $L_{y}$ which is polynomial functions of $A_{x}$ and $B_{y}$, so that,
@@
$\forall x,x^{\prime}\ \ A_{x}A_{x^{\prime}}+A_{x}A_{x^{\prime}}=2\delta_{x,x^{\prime}}\mathbb{I},$ (40)
This relation is satisfied when the operators $A_{x}$’s are the generators of a Clifford algebra [60]. It is known that in any dimension $d$ a maximum of $d^{2}-1$ mutually anti-commuting dichotomic observables exist [60]. Then for any value of $N$ there exists an upper bound on the dimension of the quantum system.
@@
We note that the necessary condition for $N$-wise compatibility for unsharp POVMs derived above matches the necessary condition for $N$-wise compatibility of the same class of POVMs derived in [2, 3]. However, while deriving the necessary condition they have used operator space structure of quantum theory itself. In contrast, staring from our theory-independent necessary condition derived in Eq.(29), we reproduced their necessary condition by inserting the operator formalism of quantum theory while optimizing the functional $(\mathcal{B}_{N})_{Q}$ in Eq. (31).
Thus, our inequality can be cast as a witness for incompatibility whose violation quantifies the amount of incompatibility a set of observables possess. It is already shown that for a set of mutually anti-commuting observables the derived inequality can be violated maximally. Such a set forms the maximal incompatible set of observables in quantum theory, as already established in literature [61]. However, in any other no-signalling theory there may not exists observables equivalent to something like a mutually anti-commuting set. For such theories other different values of unsharpness parameter $\eta^{\ast}$ (characteristics of that particular theory) is necessary to satisfy the constructed inequality for any arbitrary $N$, thereby possessing different degree of incompatibility as per the Definition 3.
@@ VII Geometric implications
In literature [40] the simplicial GPTs are referred to as the classical due to the following reason. If the extremal states are the representative of comprehensively distinct states of the system then any legitimate pure state should correspond to those extremal states with an extra proviso that any legitimate mixed state must portray the epistemic uncertainty about the extremal states. Thus every mixed state should be uniquely determined by the convex decomposition of the pure states. In discrete geometry, simplex is the only regular geometrical structure that has this specific property that any point inside the simplex can be uniquely decomposed into the points representing its vertices. However, there is no reason to think that every classical GPT would be simplicial. Nevertheless there are GPTs that are not simplicial, but they can embadded into a simplicial GPT. Such GPTs are defined as as follows.
@@
Thus there may exist GPTs that are simplex-embeddable but not simplicial. Such GPTs can then be treated as classically simulable rather than classical as noted in [62]. Moreover, it is argued in [20] that a simplex-embeddable GPT that respects no-restriction hypothesis [46] must be a simplicial GPT. In this paper however, we do not restrict ourselves only to the GPTs that respects the no-restriction hypothesis and consider the notion of simplicticity and simplex-embeddability to be two distinct layer of classicality.
Figure 3: Schematic diagram of state and effect spaces represented in three dimensional Euclidean space. While the left side of the arrow depicts an arbitrary shaped state and effects, the right side shows that the state and effect space transformation in order to embedding them into a simplex and hypercube respectively. The no-fill arrows show the result obtained in [66, 65] that suitable noise in state viable of simplex-embeddability, while the gray arrow shows the result obtained in this paper namely the unsharpness in measurement is another viable parameter of simplex-embeddability.
Now, we discuss the geometrical properties of the GPT describing Bob’s local statistics obtained from the remotely prepared systems discussed in the preceding sections in terms of its connections with the Alice’s compatible measurements. For this, we recall the result of ref. [20] that the ontological model of an operational theory is generalised noncontextual iff the GPT describing the operational statistics is simplex-embeddable. It is already clear from previous section that in a bipartite scenario with a nontrivially correlated shared state, if Alice’s measurements are all compatible then Bob’s statistics can be reproduced by a generalised noncontextual model. By connecting the result of ref. [20] and the results obtained in this paper we prove the following theorem.
@@
Proof: Consider the state space of a GPT is a unit $d$-simplex $\mathcal{S}$ with the effect space a unit hypercube $\mathcal{S}^{*}$ dual to the simplex lying in $\mathbb{R}^{d}$. This then forms a simplicial GPT according to Definition 4. Now, the proof is very straightforward from ref. [20] when one identifies the $\lambda\in\Lambda$ as the ontic states describing the steered preparations. Due to the positivity of points $\{s_{i}\}$ one easily identify the vertices of the simplex $\mathcal{S}$ is isomorphic with each of the variable $\lambda\in\Lambda$ and so with the elements of the set $a=\{a_{x}\}_{x=1}^{N}$. For instance, we consider the set $a$ as points on a vector space $\mathbb{R}^{\Lambda}$ forming $n(a)$ number of vertices of the simplex $\mathcal{S}$, with $n(a)$ being the cardinality of the set $a$.

### v3

@@ III.2 Ontological model of a GPT and generalised noncontextuality
At this point, we recall an important result from Ref. spekkens21 that will be used to establish our main result. We encapsulate that result in the following theorem and provide a straightforward proof of it.
@@ IV Notion of measurement incompatibility in a GPT
Given the state space $\Omega$, any measurement in the GPT formalism is represented by an affine map $M:\Omega\rightarrow\mathcal{P}(\Omega)$, where $\mathcal{P}(\Omega)$ constitutes the set of probabilities for the outcomes of the measurement. Technically, a measurement $M$ on a system state $\omega\in\Omega$ producinges dichotomic outcomes $m\in\{\pm 1\}$ can be represented as, $M=\{e_{\tiny M,\omega}^{\tiny m}|\sum_{m}e_{\tiny M,\omega}^{\tiny m}=1,e_{\tiny M,\omega}^{\tiny m}\geq 0,\forall\omega\in\Omega\}$. Therefore, the expected average value of $M$ can be written as
@@
Such joint effects corresponding to a pair of observables may not always exist when only perfect measurements are considered. However, for the imperfect or fuzzy measurement of a pair of observables, such a joint effect may exist if the degree of fuzziness is upper bounded by a certain critical value busch13 . We call such fuzzy version of the measurements as unsharp measurement and represent the unsharpness parameter as $\eta\in[0,1]$.
@@
$N^{\eta^{*}}$-compatible GPT: A specific GPT is said to be $N^{\eta^{*}}$-compatible GPT if the necessary condition for the $N$-wise compatibility of the fuzzy measurement of $N$ arbitrary mutually-maximally incompatible observables of that GPT is $\eta\leq\eta^{*}$.
Note that for any given value of $N$, the higher the value of $\eta^{*}$, the less $N$-incompatible the GPT is. For instance, the extreme case of $\eta^{*}=1$ indicates that the GPT is $N-$wise compatible, even if the measurements are perfect. Such GPTs are deemed classical in terms of incompatibility, as those $N$ observables have simultaneous values. Following definition 3, a GPT can be said to be maximally incompatible if in such a GPT, the highest amount of unsharpness $(\eta^{\ast})$ needs to be given to make any arbitrary $N$ observable compatible. It is shown in heinosaari16 that if $\eta^{*}=\frac{1}{N}$ then all $N$ observables are compatible in any arbitrary GPT. Therefore, a GPT is referred to as maximally incompatible if $\eta^{*}=\frac{1}{N}$ heinosaari16 . However, it may be possible that in a specific GPT for $\eta>\frac{1}{N}$ is enough to establish $N$-wise compatibility.
We note here that in quantum theory, the maximally incompatible pair of observables do not exist heinosaari16 ; heinosaari14 in finite-dimensional Hilbert space. However, if the dimension of the quantum system is infinite, one needs the unsharpness parameter $\eta=1/2$ to make an arbitrary pair of observables compatible. On the other hand, any pair of qubit observables becomes compatible if and only if the unsharpness parameter is $\eta\leq 1/\sqrt{2}$ busch . Later, this result is generalised banik13 for any pair of dichotomic observables in an arbitrary finite dimension of Hilbert space. We remark here that a very few results regarding $N$-wise compatibility in any arbitrary dimension $d$ are known. The best known wiseman07 tight bound on the unsharpness for the compatibility of set of $N$ noisy measurements is $\eta\leq\frac{H_{d}-1}{d-1}$, where , $H_{d}=\sum_{n=1}^{d}\frac{1}{n}$ is called the $d^{th}$ harmonics. It has also been argued in wiseman07 ; quintino14 ; uola14 that for any dimension $d$, the above bound is necessary and sufficient for the compatibility of arbitrary number of measurements. For example, an arbitrary $N$ noisy qubit measurements becomes compatible iff $\eta\leq 1/2$ quintino14 . By noting the connection between compatibility of a set of measurements with quantum steering, this bound is also supported by recent results chitambar24 ; renner24 .
Given the above perspective, we ask the more general question of what is the necessary condition for $N$-wise compatibility for any no-signalling GPT. From the quantum mechanical perspective, such a $N$-wise compatibility condition enables us to find the critical value of $\eta$ required to make an arbitrary $N$ number of dichotomic observables compatible.
@@ V Landscape of the arguments and Results
where $\omega^{\lambda}$ is the GPT state corresponding to variable $\lambda$. Non-existence of such decomposition of the states is usually referred to as steering schrodinger ; wiseman07 in literature. However, we do not intend to discuss the very details of steering. Instead, we use the proof for the existence of states in the form Eq. (19) as a helping tool for proving one of our main results articulated in Theorem 2, appearing later in this section.
To show that a decomposition as Eq. (19) exists for the reduced states at Bob’s wing, let us first assume that there is a one-to-one correspondence between the elements of $\Lambda$ and $a$ such that $\lambda=a_{x}$ specifies one of the remotely prepared states. The variable $\lambda$ is distributed according to the probability distribution $p(\lambda)=\langle e^{a_{x}}_{A},\omega_{A}^{u_{B}}\rangle$, where $\omega_{A}^{u_{B}}$ is Alice’s part of the state when Bob measures the unit effect $u_{B}$ on his part of the subsystem. In each run of experiment Alice sends her local state $\omega^{a_{x}}_{B}=[(e^{a_{x}}\otimes u_{B}),\omega_{AB}]$ to Bob. Furthermore, when Bob asks her to perform a specific measurement, say $x$, and announce the outcome, she simply conveys the outcome $a_{x}$ according to the probability distribution $p(a_{x}|x,\lambda)=\delta_{\lambda,a_{x}}$. Thus each of the states that Bob gets can be written as Eq. (19). The correspondence between the existence of such decomposition of remotely prepared states and the joint measurability has also been demonstrated in ref. jencova22 ; manik15j .
@@
Figure 1: Pictorial depiction of main results. Dark red arrows are findings of the current paper while the blur arrow indicates the result of ref. spekkens21 .
@@
$\mathcal{B}_{N}=\max_{B_{y}}\Big{(}\sum_{y=1}^{2^{N-1}}\lvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle-\Delta(N)\rvert\Big{)}\leq 2^{N-1}$ (29)
where $l_{y}^{x}$ is the $x^{th}$ bit sampled from $y^{th}$ string from the bit-string $l\in\{0,1\}^{n}$ having first entry 00 and $\Delta(N)$ is a function of $N$-wise joint probabilities.
@@ VI Optimal measurement incompatibility allowed in quantum theory
Consider that Alice’s measurements are described by unbiased POVMs (smeared versions of projective measurement) of the form $A_{x}\equiv\Big{\{}A_{\pm|x}\ \big{|}\ E_{\pm|x}=\frac{1}{2}(\mathbb{I}\pm\eta A_{x})\Big{\}}$, where $\eta\in[0,1]$ is the unsharpness parameter. Hence, effective observable of Alice is $\eta A_{x}$. We note that for mutually-maximally incompatible quantum observables $\Delta(N)=0$. Thus, if we assume the same unsharpness parameter $\eta$ for each of the observable $A_{x}$, we can then write the quantum value of the functional $\mathcal{B}_{N}$ as,
$(\mathcal{B}_{N})_{Q}=\eta\sum_{y=1}^{2^{N-1}}\rvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert.$ (31)
To optimize $(\mathcal{B}_{N})_{Q}$ we use an elegant sum-of-square approach pan20 as explicitly described in the following. This implies showing $(\mathcal{B}_{n})_{Q}\leq\beta_{N}$ for all possible quantum states $\rho_{AB}$ and measurement observables $A_{x}$ and $B_{y}$, where $\beta_{N}$ is the upper bound. This is equivalent to showing that there is a positive semidefinite operator $\gamma_{N}\geq 0$, that can be expressed as $\langle\gamma_{N}\rangle_{Q}=\beta_{N}-(\mathcal{B}_{n})_{Q}$. We prove this by considering a set of suitable positive operators $L_{y}$ which is polynomial functions of $A_{x}$ and $B_{y}$, so that,
@@
$\forall x,x^{\prime},\ \ A_{x}A_{x^{\prime}}+A_{x^{\prime}}A_{x}=2\delta_{x,x^{\prime}}\mathbb{I},$ (40)
This relation is satisfied when the operators $A_{x}$’s are the generators of a Clifford algebra clifford . It is known that in any dimension $d$ a maximum of $d^{2}-1$ mutually anti-commuting dichotomic observables exist clifford . Then for any value of $N$ there exists an upper bound on the dimension of the quantum system.
@@
We note that the necessary condition for $N$-wise compatibility for unsharp POVMs derived above matches the necessary condition for $N$-wise compatibility of the same class of POVMs derived in kunjwal14 ; kunjwal14g . However, while deriving the necessary condition they have used operator space structure of quantum theory itself. In contrast, staring from our theory-independent necessary condition derived in Eq.(29), we reproduced their necessary condition by inserting the operator formalism of quantum theory while optimizing the functional $(\mathcal{B}_{N})_{Q}$ in Eq. (31).
Thus, our inequality can be cast as a witness for incompatibility whose violation quantifies the amount of incompatibility a set of observables possess. It is already shown that for a set of mutually anti-commuting observables the derived inequality can be violated maximally. Such a set forms the maximal incompatible set of observables in quantum theory, as already established in literature winter08 . However, in any other no-signalling theory there may not exists observables equivalent to something like a mutually anti-commuting set. For such theories other different values of unsharpness parameter $\eta^{\ast}$ (characteristics of that particular theory) is necessary to satisfy the constructed inequality for any arbitrary $N$, thereby possessing different degree of incompatibility as per the Definition 3.
@@ VII Geometric implications
In literature barrett07 the simplicial GPTs are referred to as the classical due to the following reason. If the extremal states are the representative of comprehensively distinct states of the system then any legitimate pure state should correspond to those extremal states with an extra proviso that any legitimate mixed state must portray the epistemic uncertainty about the extremal states. Thus every mixed state should be uniquely determined by the convex decomposition of the pure states. In discrete geometry, simplex is the only regular geometrical structure that has this specific property that any point inside the simplex can be uniquely decomposed into the points representing its vertices. However, there is no reason to think that every classical GPT would be simplicial. Nevertheless there are GPTs that are not simplicial, but they can embadded into a simplicial GPT. Such GPTs are defined as as follows.
@@
Thus there may exist GPTs that are simplex-embeddable but not simplicial. Such GPTs can then be treated as classically simulable rather than classical as noted in shahandeh21 . Moreover, it is argued in spekkens21 that a simplex-embeddable GPT that respects no-restriction hypothesis janotta13 must be a simplicial GPT. In this paper however, we do not restrict ourselves only to the GPTs that respects the no-restriction hypothesis and consider the notion of simplicticity and simplex-embeddability to be two distinct layer of classicality.
Figure 3: Schematic diagram of state and effect spaces represented in three dimensional Euclidean space. While the left side of the arrow depicts an arbitrary shaped state and effects, the right side shows that the state and effect space transformation in order to embedding them into a simplex and hypercube respectively. The no-fill arrows show the result obtained in rossi23 ; selby24 that suitable noise in state viable of simplex-embeddability, while the gray arrow shows the result obtained in this paper namely the unsharpness in measurement is another viable parameter of simplex-embeddability.
Now, we discuss the geometrical properties of the GPT describing Bob’s local statistics obtained from the remotely prepared systems discussed in the preceding sections in terms of its connections with the Alice’s compatible measurements. For this, we recall the result of ref. spekkens21 that the ontological model of an operational theory is generalised noncontextual iff the GPT describing the operational statistics is simplex-embeddable. It is already clear from previous section that in a bipartite scenario with a nontrivially correlated shared state, if Alice’s measurements are all compatible then Bob’s statistics can be reproduced by a generalised noncontextual model. By connecting the result of ref. spekkens21 and the results obtained in this paper we prove the following theorem.
@@
Proof: Consider the state space of a GPT is a unit $d$-simplex $\mathcal{S}$ with the effect space a unit hypercube $\mathcal{S}^{*}$ dual to the simplex lying in $\mathbb{R}^{d}$. This then forms a simplicial GPT according to Definition 4. Now, the proof is very straightforward from ref. spekkens21 when one identifies the $\lambda\in\Lambda$ as the ontic states describing the steered preparations. Due to the positivity of points $\{s_{i}\}$ one easily identify the vertices of the simplex $\mathcal{S}$ is isomorphic with each of the variable $\lambda\in\Lambda$ and so with the elements of the set $a=\{a_{x}\}_{x=1}^{N}$. For instance, we consider the set $a$ as points on a vector space $\mathbb{R}^{\Lambda}$ forming $n(a)$ number of vertices of the simplex $\mathcal{S}$, with $n(a)$ being the cardinality of the set $a$.